# Minimizing the number of edges in LC-equivalent graph states

### v1

@@ Abstract
Graph states are a powerful class of entangled states with numerous applications in quantum communication and quantum computation. Local Clifford (LC) operations that map one graph state to another can alter the structure of the corresponding graphs, including changing the number of edges. Here, we tackle the associated edge-minimization problem: finding graphs with the minimum number of edges in the LC-equivalence class of a given graph. Such graphs are called minimum edge representatives (MER), and are crucial for minimizing the resources required to create a graph state. We leverage Bouchet’s algebraic formulation of LC-equivalence to encode the edge-minimization problem as an integer linear program (ILP). We further propose a simulated annealing (SA) approach guided by the local clustering coefficient for edge minimization. We identify new MERs for graph states with up to 16 qubits by combining SA and ILP. We extend the ILP to weighted-edge minimization, where each edge has an associated weight, and prove that this problem is NP-complete. Finally, we employ our tools to minimize resources required to create all-photonic generalized repeater graph states using fusion operations.
I Introduction
Graph states are a powerful yet tractable class of multipartite entangled states that have been extensively studied in quantum science. Beyond their theoretical significance [1, 2], they support a wide range of applications, including measurement-based quantum computation [3, 4, 5], all-photonic quantum repeaters [6, 7, 8, 9, 10, 11], loss-tolerant error-correcting codes [12, 13, 14], and quantum metrology [15]. Significant efforts have been directed toward both their experimental realization [16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and theoretical optimization, encompassing photonic graph state generation [26, 27, 28, 29, 30], general graph state construction [31], and state distribution [32].
A graph state |$\ket{G}$ is represented by a graph $G$ with vertices that correspond to qubits initialized in the $\ket{+}$ state, and edges that correspond to controlled-Z (CZ) gates applied to pairs of qubits. Graph states can also be defined by Pauli stabilizers, and thus constitute a subset of the stabilizer states. By applying local Clifford (LC) gates, i.e. single-qubit Clifford unitaries, one can transform a graph state |$\ket{G}$ into a different local Clifford equivalent (or LC-equivalent) graph state |$\ket{H}$ with possibly fewer edges in its graph. It was shown in Ref. [33] that two graph states are LC-equivalent if and only if their graphs are related by a sequence of graph operations called local complementations.
@@ I Introduction
In this work, we introduce three algorithms to address the edge-minimization problem: given a graph $G$, find an LC-equivalent graph $H$ with the minimum number of edges. Our first algorithm, SA, uses simulated annealing [35] to obtain approximate solutions for graphs with up to 100 qubits. The second algorithm, ILP, is an integer linear program based on Bouchet’s algebraic characterization of LC-equivalence [36]. While ILP guarantees globally optimal solutions, its runtime scales exponentially in the worst case. Combining both approaches, our third method, SA+ILP, uses SA to precondition the input to ILP, yielding a significant constant-factor speedup. This hybrid approach enables us to compute exact MERs for graphs with up to 16 qubits. Moreover, we extend ILP to tackle the weighted-edge minimization problem: finding LC-equivalent graphs that minimize the total edge weight. We prove this problem is NP-complete via a reduction from the vertex-minor problem [37].
@@
To demonstrate the practical utility of our approach, we use SA to optimize the resources required for the creation of all-photonic generalized repeater graph states (gRGS) [11]. These states can be utilized for sharing multiple units of entanglement in a quantum network. Moreover, they can be constructed using single-photon sources and fusion operations [38]. However, the probabilistic nature of fusion and photon loss leads to a high number of required single photons and fusion operations, with the fusion order critically affecting generation efficiency. To reduce resource requirements, we identify a subgraph of the gRGS and construct its corresponding MER, which contains fewer edges and is therefore more efficient to create. LC gates are applied at the start of the protocol to transform the MER into the desired subgraph, allowing for an optimized fusion order. Using the tools from Ref. [28], we quantify the resources needed and determine the optimal fusion sequence. Our method achieves more than an order-of-magnitude reduction in resource usage under realistic high-loss conditions.
@@
In Section II, we provide a description of SA and ILP algorithms we use to find MERs. In Section III, we study the performance of our algorithms. Next, in Section IV, we demonstrate the applications of our framework for edge minimization. We conclude with an outlook for future work in Section V.
II Algorithms for edge minimization
II.1 Simulated annealing
@@
Here we discuss a heuristic approach for edge minimization based on simulated annealing (SA) [35]. Due to its flexibility, simulated annealing has been widely applied in various domains, including the distribution of graph states in networks [32].
@@
SA is an iterative algorithm that runs for a given number of iterations ($k_{\textrm{max}}$) while trying to approximate the optimal solution. At each iteration $k$, SA may move from a state $s_{k}$ to another potential ‘neighboring’ state $s_{\textrm{pot}}$ with a certain acceptance probability. In our implementation, states correspond to graphs, and the state $s_{\textrm{pot}}$ is sampled from a set of neighboring states $S(s_{k})$. Here, $S(s_{k})$ denotes the subset of graph states that differ from the graph state $s_{k}$ by one local complementation. A local complementation $L_{v}$ is an operation on a graph $G$ that acts on a vertex $v$ by complementing the edges in the neighborhood of $v$, as shown in Fig. 1. We construct the set $S(s_{k})$ based on a metric called the local clustering coefficient [39], and the degree $D_{v}$ of a vertex $v$. Details for finding $S(s_{k})$ are discussed in Appendix C.
@@
Figure 1: Local complementation on the red vertex removes (or adds) an edge between the neighbors of the vertex if they were connected (or disconnected) prior to its application
For efficient SA, the temperature schedule, i.e., the temperature $T(k)$ at each iteration $k$, must be properly defined. We consider a logarithmic cooling schedule [40]: at the $k$’th iteration, the temperature is given by $T(k)=T(1)/\log_{2}(k+1)$, for a given initial temperature $T(1)$. In principle, the initial temperature can be tailored for specific problem instances; however, we found $T(1)=50$ and $T(1)=100$ to be good enough in practice to find approximate MERs.
II.2 Edge minimization using Bouchet’s algorithm
We leverage Bouchet’s LC-equivalence between graph states to find the exact solutions to the edge-minimization problem. By definition, the edge-minimization problem can be written as follows:
@@ II.2 Edge minimization using Bouchet’s algorithm
where $A_{G}~{}(A_{H})$ represents the adjacency matrix of $G~{}(H)$. We use Bouchet’s algorithm for LC-equivalence [36] to express LC-equivalence in a form amenable to (integer) linear programming. For completeness, we describe Bouchet’s algorithm in Appendix A. The algorithm takes as input the adjacency matrices $A_{G}$ and $A_{H}$ of $n$-vertex graphs $G$ and $H$ respectively, and is based on the binary/symplectic framework (see for example [41, 42] and Appendix A for further information) to find the following equation describing LC-equivalence of graphs $G$ and $H$,
@@
We then linearize the equations by introducing constraints to convert products of integer variables into linear systems. This is done for products of variables originating from the terms $A_{H}P$ and $A_{H}R$ in Eq. (4) and terms $PS$ and $RQ$ in Eq. (2). The approach outlined above produces the full set of constraints; since these are rather unwieldy, we defer the complete derivation of the ILP algorithm to Appendix B.
@@
The runtime of the ILP algorithm depends on the number of constraints, which in turn depend on the number of vertices and edges in the input graphs. Therefore, we employ SA to find the approximate MER of a given graph before inputting it into ILP. This preprocessing step reduces the number of constraints in the corresponding ILP thus improving the runtime. We refer to this combined algorithm as SA+ILP. We use the MOSEK optimizer API for Python [43] to run ILP and SA+ILP. Since ILP is an exact optimization, the results obtained are global minima, i.e. MERs of their corresponding LC-orbit. In Section III, we benchmark the performance of our algorithms for different graph models.
@@
Figure 2: In (a) and (c), we compare the number of edges in the input graphs and the outputs from SA and SA+ILP for Erdős-Rényi graphs and bounded-degree graphs respectively. The solid lines denote the average value, and the shaded bands indicate the maximum and minimum. (b) shows the runtime of the SA+ILP algorithm as a function of edge inclusion probability for ER graphs with 13 vertices. (d) compares the runtime between ILP and SA+ILP for bounded-degree graphs. The solid line indicates the average runtime, and error bars show the 95th percentile around the mean. For ER graphs, the runtime and the MER edge count are highest around $p\sim 0.5$. From (d), we observe that SA+ILP achieves a constant-factor reduction in runtime compared to ILP on average. This improvement can be attributed to the SA preprocessing step, although the runtime remains exponential in the size of the input graph. The average runtime of SA+ILP for graphs with 16 vertices is smaller than that for graphs with 15 vertices, which we attribute to a better SA performance.
II.3 Weighted-edge minimization
@@
Our formulation of Bouchet’s algorithm as an optimization problem can be easily extended to include weighted-edge minimization of LC-equivalent graph states. The optimization problem then becomes
@@
where $W$ is a weighted adjacency matrix encoding the weights $W[i,j]$ of each edge $(i,j)$. These weights could represent, for example, the cost of creating edges, thereby enabling a more realistic analysis of the resource requirements for generating graph states, similar to the simulated annealing based approach of Ref. [32].
The decision problem for arbitrary weights can be proved to be NP-complete by reducing the so-called vertex-minor problem to the weighted edge-minimization problem. The vertex-minor (or qubit-minor) problem plays an important role in quantum information theory [44, 45]. A graph state |$\ket{H}$ that can be obtained from |$\ket{G}$ under local Clifford unitaries and Pauli measurements is called a qubit-minor of |$\ket{G}$ [44, 45]. Accordingly, the graph $H$ is called a vertex-minor of graph $G$ [46]. We consider the labeled version of the vertex-minor problem, where the resultant graph after measurements and local complementations needs to be exactly $H$ (not up to graph isomorphism). While local equivalence of two graphs $G$ and $H$ can be decided efficiently, deciding whether $H$ is a vertex-minor of $G$ is NP-complete, for both the labeled [45] and unlabeled case [37].
The vertex-minor problem can be formulated using the weighted-edge minimization problem as follows. Assume that the graph $H=(V^{\prime},E^{\prime})$ has no isolated vertices  [^1] For the case of $H$ containing isolated vertices, assign a weight of 00 to any of the edges incident to an isolated vertex.. We define the weights $W[i,j]$ according to the following rule:
@@
The smallest weight possible is -$\left|E^{\prime}\right|$, which is achievable if and only if $G$ is LC-equivalent to a graph $\overline{G}$ such that $\overline{G}\left[V^{\prime}\right]=H$, i.e. when $H$ is a vertex-minor of $G$. In other words, an efficient algorithm for deciding whether the minimum score equals $-\left|E^{\prime}\right|$ would also yield an efficient method for determining whether $H$ is a vertex-minor of $G$, and vice versa. Since the vertex-minor problem is known to be NP-complete in general [37], it follows immediately that the decision problem for weighted-edge minimization is NP-complete as well. We note that this still leaves open whether the minimization problem remains NP-complete when restricting to uniform/non-negative weights.
@@
Figure 3: (a) demonstrates the fusion order for creating an RGS with eight qubits, obtained using OptGraphState from Ref. [28]. The order corresponds to the graph $H$+$L$ with LC-gates applied at the beginning. Graph states (shown with blue vertices and solid edges) may transform into non-graph states when acted upon by LC-gates (represented with red and yellow vertices). The resulting state might also be a non-graph state if such a qubit is involved in fusions, indicated using dashed edges. (b) shows: (1) the sampled gRGS state capable of sharing three ebits across a quantum network, (2) the central graph $G$ (with 46 edges), and (3) its approximate MER $H$, which contains 23 edges.
III Edge minimization of LC-equivalent graph states
@@
In this section, we first characterize the runtime of SA. We then empirically evaluate: (1) the runtime of ILP and SA+ILP, and (2) the reduction in edge count achieved by SA and SA+ILP. We focus on two graph models: (1) the Erdős-Rényi model, which is a standard model for creating random graphs, and (2) the bounded-degree model, which encompasses realistic hardware constraints. Our algorithms ran on each input graph by allocating them on 20 cores of Intel(R) Xeon(R) CPU E5-2620 v2 @ 2.10GHz, each with 32GB of memory.
@@
We first benchmark our algorithms for the Erdős-Rényi (ER) graph model [48, 49]. An ER graph $G(n,p)$ is a random graph on $n$ vertices, where each of the $\frac{n(n-1)}{2}$ possible edges is included independently with a probability $p$. The expected number of edges in the initial graph increases as the probability $p$ increases. We can thus control the sparsity (or density) of edges in the generated graphs by changing the value of $p$. We first characterize the runtime of the SA algorithm by running it for different numbers of maximum iterations. The initial temperature is set to $T(1)=50$ for a set of 100 Erdős-Rényi graphs with 100 vertices and 2971.2 edges on average. For $k_{\textrm{max}}=50$, SA had an average runtime of 0.94 seconds, and the number of edges obtained in the output graphs was 2367.26 on average. Increasing $k_{\textrm{max}}$ to 1050 increases the runtime to 18.82 seconds, and the number of edges obtained was 2310.73. In the rest of the benchmarking, we set the value of $k_{\textrm{max}}$ and $T(1)$ to 100.
@@
We generate graphs for increasing values of $p$ with $n=13$ vertices using the ER graph model. We generate 100 ER graphs for each value of $p$, and we apply SA and SA+ILP to minimize the edges in each graph generated for a $p$ value. From Fig. 2(a), we find that the number of edges obtained from SA is on average 1.07 times larger than the MERs for ER graphs. The expected number of edges is concave in $p$, with a maximum at around $p\sim\frac{1}{2}$, and SA approximations are slightly worse near this regime. Moreover, we observe that dense graphs ($p\geq 0.6$) are LC-equivalent to graphs with relatively fewer edges, shown by the large gap between the input edge count and MER edge count. In Fig. 2(b), the correlation coefficient between the logarithm of runtime of SA+ILP and the number of input edges is 0.94, confirming an exponential runtime dependence on the size of input graph. Thus, we omit running the ILP algorithm for ER graphs in Fig. 2(b) since the number of edges grows to a large number ($\sim 70$) for higher values of probability $p$.
@@
We also benchmark our algorithms using the bounded-degree graph model, which incorporates hardware constraints that caps the maximum degree of edges like restricted qubit connectivity. We generate random graphs with $n$ vertices and bounded degree of $d_{\textrm{lim}}$ by sampling a set of $n$ positive integers where all elements are $\leq d_{\textrm{lim}}$ and constructing simple connected graphs using the Havel-Hakimi algorithm [50, 51]. In our simulations, we fixed the value of $d_{\textrm{lim}}$ to 5 and generated graphs with $n$ ranging from 6 to 16. For each vertex number, we generated 100 graphs; however, only 63 such graphs could be generated for $n=6$ vertices.
@@
In Fig. 2(c), we benchmark the edge minimization performance of SA against the MERs obtained from SA+ILP for bounded-degree graphs. The number of edges obtained from SA is on average 1.04 times larger than the MERs. We observe that the performance of SA declines slightly as the number of vertices increases, likely due to the increased sparsity of the input graphs. In addition to edge count comparison, we compare the runtime of ILP and SA+ILP in Fig. 2(d). To find the dependence of runtime of SA+ILP, we calculate the correlation coefficient between the logarithm of its runtime and the number of edges and vertices input to the ILP part of the algorithm. The correlation coefficients between the logarithm of runtime and input edges (vertices) were 0.96 (0.89). Moreover, the addition of SA as a pre-processing step to ILP reduces the runtime of the SA+ILP algorithm by a factor of two compared to ILP alone for bounded-degree graphs. Thus, SA+ILP is a faster algorithm than ILP but it still has an exponential runtime dependence on the number of edges and vertices in the input graphs.
IV Creation of generalized repeater graph states

### v2

# Minimising the number of edges in LC-equivalent graph states

@@ Abstract
Graph states are a powerful class of entangled states with numerous applications in quantum communication and quantum computation. Local Clifford (LC) operations that map one graph state to another can alter the structure of the corresponding graphs, including changing the number of edges. Here, we tackle the associated edge-minimisation problem: finding graphs with the minimum number of edges in the LC-equivalence class of a given graph. Such graphs are called minimum edge representatives (MER) and are crucial for minimising the resources required to create a graph state. We leverage Bouchet’s algebraic formulation of LC-equivalence to encode the edge-minimisation problem as an integer linear program (EDM-ILP). We further propose a simulated annealing (EDM-SA) approach guided by the local clustering coefficient for edge minimisation. We identify new MERs for graph states with up to 16 qubits by combining EDM-SA and EDM-ILP. We extend the ILP to weighted-edge minimisation, where each edge has an associated weight, and prove that this problem is NP-complete. Finally, we employ our tools to minimise the resources required to create all-photonic generalised repeater graph states using fusion operations.
1 Introduction
Graph states are a powerful yet tractable class of multipartite entangled states that have been extensively studied in quantum science. Beyond their theoretical significance [1, 2], they support a wide range of applications, including measurement-based quantum computation [3, 4, 5, 6, 7, 8], all-photonic quantum repeaters [9, 10, 11, 12, 13, 14], loss-tolerant error-correcting codes [15, 16, 17], and quantum metrology [18]. Significant efforts have been directed toward both their experimental realisation [19, 20, 21, 22, 23, 24, 25, 26, 27, 28] and theoretical optimisation, encompassing photonic graph state generation [29, 30, 31, 32, 33], general graph state construction [34], and state distribution [35].
[…]
@@ 1 Introduction
In this work, we introduce three algorithms to address the edge-minimisation problem: given a graph $G$, find an LC-equivalent graph $H$ with the minimum number of edges. Our first algorithm, EDM-SA, uses simulated annealing [38] to obtain approximate solutions for graphs with up to 100 qubits. The second algorithm, EDM-ILP, is an integer linear program based on Bouchet’s algebraic characterisation of LC-equivalence [39]. While EDM-ILP guarantees globally optimal solutions, its runtime scales exponentially in the worst case. Combining both approaches, our third method, EDM-SAILP, uses EDM-SA to precondition the input to EDM-ILP, yielding a significant constant-factor speed-up. This hybrid approach enables us to compute exact MERs for graphs with up to 16 qubits. Moreover, we extend EDM-ILP to tackle the weighted-edge minimisation problem: finding LC-equivalent graphs that minimise the total edge weight. We prove this problem is NP-complete via a reduction from the vertex-minor problem [40].
@@
To demonstrate the practical utility of our approach, we use EDM-SA to optimise the resources required for the creation of all-photonic generalised repeater graph states (gRGS) [14]. These states can be utilised for sharing multiple units of entanglement in a quantum network. Moreover, they can be constructed using single-photon sources and fusion operations [41]. However, the probabilistic nature of fusion and photon loss leads to a high number of required single photons and fusion operations, with the fusion order critically affecting generation efficiency. To reduce resource requirements, we identify a sub-graph of the gRGS and construct its corresponding MER, which contains fewer edges and is therefore more efficient to create. LC gates are applied at the start of the protocol to transform the MER into the desired sub-graph, allowing for an optimised fusion order. Using the tools from Ref. [31], we quantify the resources needed and determine the optimal fusion sequence. Our method achieves more than an order-of-magnitude reduction in resource usage under realistic high-loss conditions.
@@
In Section 2, we provide a description of EDM-SA and EDM-ILP algorithms we use to find MERs. In Section 3, we study the performance of our algorithms. Next, in Section 4, we demonstrate the applications of our framework for edge minimisation. We conclude with an outlook for future work in Section 5.
2 Algorithms for edge minimisation
2.1 Simulated annealing
@@
Here we discuss a heuristic approach for edge minimisation based on simulated annealing (SA) [38]. Due to its flexibility, simulated annealing has been widely applied in various domains, including the optimisation of resource states in measurement-based quantum computation [6, 8], and the distribution of graph states in networks [35].
@@
SA is an iterative algorithm that runs for a given number of iterations ($k_{\textrm{max}}$) while trying to approximate the optimal solution. At each iteration $k$, SA may move from a state $s_{k}$ to another potential ‘neighbouring’ state $s_{\textrm{pot}}$ with a certain acceptance probability. In our implementation, states correspond to graphs, and the state $s_{\textrm{pot}}$ is sampled from a set of neighbouring states $S(s_{k})$. Here, $S(s_{k})$ denotes the subset of graph states that differ from the graph state $s_{k}$ by one local complementation. A local complementation $L_{v}$ is an operation on a graph $G$ that acts on a vertex $v$ by complementing the edges in the neighbourhood of $v$, as shown in Fig. 1. We construct the set $S(s_{k})$ based on a metric called the local clustering coefficient [42], and the degree $D_{v}$ of a vertex $v$. Details for finding $S(s_{k})$ are discussed in Appendix C.
@@
Figure 1: Local complementation on the red vertex removes (or adds) an edge between the neighbours of the vertex if they were connected (or disconnected) prior to its application
For efficient SA, the temperature schedule, i.e., the temperature $T(k)$ at each iteration $k$, must be properly defined. We consider a logarithmic cooling schedule [43]: at the $k$’th iteration, the temperature is given by $T(k)=T(1)/\log_{2}(k+1)$, for a given initial temperature $T(1)$. Logarithmic cooling schedule is slower compared to other cooling schedules. This allows EDM-SA algorithm to explore the states for longer time before converging to a solution. In principle, the initial temperature and $k_{\textrm{max}}$ can be tailored for specific problem instances. However, we found $T(1)=50$ and $T(1)=100$ to be good enough in practice to find approximate MERs.
2.2 Edge minimisation using Bouchet’s algorithm
We leverage Bouchet’s LC-equivalence between graph states to find the exact solutions to the edge-minimisation problem. By definition, it can be written as follows:
@@ 2.2 Edge minimisation using Bouchet’s algorithm
where $A_{G}~(A_{H})$ represents the adjacency matrix of G(H)G~(H). We use Bouchet’s algorithm for LC-equivalence [39] to express LC-equivalence in a form amenable to (integer) linear programming. For completeness, we describe Bouchet’s algorithm in Appendix A. The algorithm takes as input the adjacency matrices $A_{G}$ and $A_{H}$ of $n$-vertex graphs $G$ and $H$ respectively, and is based on the binary/symplectic framework (see for example [44, 45] and Appendix A for further information) to find the following equation describing LC-equivalence of graphs $G$ and $H$,
@@
We then linearise the equations by introducing constraints to convert products of integer variables into linear systems. This is done for products of variables originating from the terms $PA_{H}$ and $A_{H}R$ in Eq. (4) and terms $PS$ and $RQ$ in Eq. (2). The approach outlined above produces the full set of constraints; since these are rather unwieldy, we defer the complete derivation of the EDM-ILP algorithm to Appendix B. We note that Eq. (1) form $n^{2}$ equations in terms of $4n$ variables, but with the non-linear constraint in Eq. (2). Bouchet showed in [39] that one only needs to check a small number of solutions, to certify whether there exists a solution that also satisfies Eq. (2). In particular, if $U$ forms any basis of the vector space $\mathcal{V}$ of solutions to Eq. (1), then there exists a solution in $\mathcal{V}$ that satisfies the constraint in (4) if and only if there exists an element in the set $\{{u+u^{\prime}\mid u,u^{\prime}\in U\}}$ that satisfies the constraint. Note that this latter set has size at most linear in $n$, while $\mathcal{V}$ can in principle be of size exponential in $n$. Unfortunately, we were not able to exploit this structure for the ILP, which has the potential to reduce the number of constraints and thus the runtime significantly. One obstacle encountered is the fact that bases are not naturally expressed in terms of ILPs, making it hard to exploit the above structure. On the other hand, given access to $U$, it is possible to express the set $\{{u+u^{\prime}\mid u,u^{\prime}\in U\}}$ naturally in terms of an ILP.
@@
The runtime of the EDM-ILP algorithm depends on the number of constraints, which in turn depend on the number of vertices and edges in the input graphs. Therefore, we employ EDM-SA to find the approximate MER of a given graph before inputting it into EDM-ILP. This preprocessing step reduces the number of constraints in the corresponding EDM-ILP thus improving the runtime. We refer to this combined algorithm as EDM-SAILP. We use the MOSEK optimiser API for Python [46] to run EDM-ILP and EDM-SAILP. Since EDM-ILP is an exact optimisation, the results obtained are global minima, i.e. MERs of their corresponding LC-orbit. In Section 3, we benchmark the performance of our algorithms for different graph models.
@@
Figure 2: In (a) and (c), we compare the number of edges in the input graphs and the outputs from EDM-SA and EDM-SAILP for Erdős-Rényi graphs and bounded-degree graphs respectively. The solid lines denote the average value, and the shaded bands indicate the maximum and minimum. (b) shows the runtime of the EDM-SAILP algorithm as a function of edge inclusion probability for ER graphs with 13 vertices. (d) compares the runtime between EDM-ILP and EDM-SAILP for bounded-degree graphs. The solid line indicates the average runtime, and error bars show the 95th percentile around the mean. For ER graphs, the runtime and the MER edge count are highest around $p\sim 0.5$. From (d), we observe that EDM-SAILP achieves a constant-factor reduction in runtime compared to EDM-ILP on average. This improvement can be attributed to the EDM-SA preprocessing step, although the runtime remains exponential in the size of the input graph. The average runtime of EDM-SAILP for graphs with 16 vertices is smaller than that for graphs with 15 vertices, which we attribute to a better EDM-SA performance.
2.3 Weighted-edge minimisation
@@
Our formulation of Bouchet’s algorithm as an optimisation problem can be easily extended to include weighted-edge minimisation of LC-equivalent graph states. The optimal solution for our case is thus the one that minimises the sum of these weights of LC-equivalent graph states. The optimisation problem then becomes
@@
where $W$ is a weighted adjacency matrix encoding the weights $W[i,j]$ of each edge possible pair $i,j$. These weights could represent, for example, the cost of creating edges, thereby enabling a more realistic analysis of the resource requirements for generating graph states, similar to the simulated annealing based approach of Ref. [35]. We allow the weights $W[i,j]$ to be real numbers, since an ILP only requires the variables to be integers, not the cost function. Note that the weights $W[i,j]$ are separate from the weights of weighted graph states, studied in e.g. [47, 48], which are not the subject of our work.
@@
The vertex-minor problem can be formulated using the weighted-edge minimisation problem as follows. Assume that the graph $H=(V^{\prime},E^{\prime})$ has no isolated vertices  [^1] For the case of $H$ containing isolated vertices, assign a weight of 0 to any of the edges incident to an isolated vertex.. We define the weights $W[i,j]$ according to the following rule:
$W[i,j]=\begin{cases*}-1&{if } $\{{i,j\}}\in E^{\prime}$\\ +1&{if } $i,j\in V^{\prime}\land\{{i,j\}}\notin E^{\prime}$\\ 0&{else}\end{cases*}\ .$ (5)
@@
The smallest weight possible is -$\left|E^{\prime}\right|$, which is achievable if and only if $G$ is LC-equivalent to a graph $\overline{G}$ such that $\overline{G}\left[V^{\prime}\right]=H$, i.e. when $H$ is a vertex-minor of $G$. In other words, an efficient algorithm for deciding whether the minimum score equals $\left|E^{\prime}\right|$ would also yield an efficient method for determining whether $H$ is a vertex-minor of $G$, and vice versa. Since the vertex-minor problem is known to be NP-complete in general [40], it follows immediately that the decision problem for weighted-edge minimisation is NP-complete as well. We note that this still leaves open whether the minimisation problem remains NP-complete when restricting to uniform/non-negative weights.
@@
Figure 3: (a) demonstrates the fusion order for creating an RGS with eight qubits, obtained using OptGraphState from Ref. [31]. The elliptical boxes show the required fusions, labelled by numbers next to them. These numbers allow the tracking of fusions after LC-gates are commuted to the beginning. The order corresponds to the graph HH+$L$ with LC-gates applied at the beginning. Graph states (shown with blue vertices and solid edges) may transform into non-graph states when acted upon by LC-gates (represented with red and yellow vertices). The resulting state might also be a non-graph state if such a qubit is involved in fusions, indicated using dashed edges. (b) shows: (1) the sampled gRGS state capable of sharing three ebits across a quantum network, (2) the central graph $G$ (with 46 edges), and (3) its approximate MER $H$, which contains 23 edges.
3 Edge minimisation of LC-equivalent graph states
@@
In this section, we first characterise the runtime of EDM-SA. We then empirically evaluate: (1) the runtime of EDM-ILP and EDM-SAILP, and (2) the reduction in edge count achieved by EDM-SA and EDM-SAILP. We focus on two graph models: (1) the Erdős-Rényi model, which is a standard model for creating random graphs, and (2) the bounded-degree model, which encompasses realistic hardware constraints. Our algorithms ran on each input graph by allocating them on 20 cores of Intel(R) Xeon(R) CPU E5-2620 v2 @ 2.10GHz, each with 32GB of memory.
@@
We first benchmark our algorithms for the Erdős-Rényi (ER) graph model [52, 53]. An ER graph $G(n,p)$ is a random graph on $n$ vertices, where each of the $\frac{n(n-1)}{2}$ possible edges is included independently with a probability $p$. The expected number of edges in the initial graph increases as the probability $p$ increases. We can thus control the sparsity (or density) of edges in the generated graphs by changing the value of $p$. We first characterise the runtime of the EDM-SA algorithm by running it for different numbers of maximum iterations. The initial temperature is set to $T(1)=50$ for a set of 100 Erdős-Rényi graphs with 100 vertices and 2971.2 edges on average. For $k_{\textrm{max}}=50$, EDM-SA had an average runtime of 0.94 seconds, and the number of edges obtained in the output graphs was 2367.26 on average. Increasing $k_{\textrm{max}}$ to 1050 increases the runtime to 18.82 seconds, and the number of edges obtained was 2310.73. This suggests that larger values of $k_{\textrm{max}}$ might allow for better performance of EDM-SA. However, in the rest of the benchmarking, we set the value of $k_{\textrm{max}}$ and $T(1)$ to 100.
@@
We generate graphs for increasing values of $p$ with $n=13$ vertices using the ER graph model. We generate 100 ER graphs for each value of $p$, and we apply EDM-SA and EDM-SAILP to minimise the edges in each graph generated for a $p$ value. From Fig. 2(a), we find that the number of edges obtained from EDM-SA is on average 1.07 times larger than the MERs for ER graphs. The expected number of edges is concave in $p$, with a maximum at around $p\sim\frac{1}{2}$, and EDM-SA approximations are slightly worse near this regime. Moreover, we observe that dense graphs ($p\geq 0.6$) are LC-equivalent to graphs with relatively fewer edges, shown by the large gap between the input edge count and MER edge count. In Fig. 2(b), the correlation coefficient between the logarithm of runtime of EDM-SAILP and the number of input edges is 0.94, confirming an exponential runtime dependence on the size of input graph. Thus, we omit running the EDM-ILP algorithm for ER graphs in Fig. 2(b) since the number of edges grows to a large number ($\sim 70$) for higher values of probability $p$.
@@
We also benchmark our algorithms using the bounded-degree graph model, which incorporates hardware constraints that caps the maximum degree of edges like restricted qubit connectivity. We generate random graphs with $n$ vertices and bounded degree of $d_{\textrm{lim}}$ by sampling a set of $n$ positive integers where all elements are $\leq d_{\textrm{lim}}$ and constructing simple connected graphs using the Havel-Hakimi algorithm [54, 55]. In our simulations, we fixed the value of $d_{\textrm{lim}}$ to 5 and generated graphs with $n$ ranging from 6 to 16. We generated 100 graphs for each vertex number; however for $n=6$ only 63 such graphs were generated.
@@
In Fig. 2(c), we benchmark the edge minimisation performance of EDM-SA against the MERs obtained from EDM-SAILP for bounded-degree graphs. The number of edges obtained from EDM-SA is on average 1.04 times larger than the MERs. We observe that the performance of EDM-SA declines slightly as the number of vertices increases, likely due to the increased sparsity of the input graphs. In addition to edge count comparison, we compare the runtime of EDM-ILP and EDM-SAILP in Fig. 2(d). To find the dependence of runtime of EDM-SAILP, we calculate the correlation coefficient between the logarithm of its runtime and the number of edges and vertices input to the EDM-ILP part of the algorithm. The correlation coefficients between the logarithm of runtime and input edges (vertices) were 0.96 (0.89). Moreover, the addition of EDM-SA as a pre-processing step to EDM-ILP reduces the runtime of the EDM-SAILP algorithm by a factor of two compared to EDM-ILP alone for bounded-degree graphs. Thus, EDM-SAILP is a faster algorithm than EDM-ILP but it still has an exponential runtime dependence on the number of edges and vertices in the input graphs.
4 Creation of generalised repeater graph states